kyles age is x, his wife is 7 years older and his son is one third of his age. the total of all their ages is 91. how old is kyle
Answer by Cromlix(4381)

Hi there,
Kyle's age = x
His wife's age = x + 7
His son's age = x/3
Total of all ages is 91.
x + (x + 7) + x/3 = 91
Taking to base 3
3x + 3(x + 7) + x = 91 x 3
3x + 3x + 21 + x = 273
Collect like terms:
3x + 3x + x = 273 - 21
7x = 252
x = 36
Kyle's age = 36 years old.
Wife's age = 43 years old.
Son's age = 12 years old.
Hope this helps :-)